obsidian1:

Which expression is an equivalent expression of 12x + 10 + 4y? A. 2(6x + 2y + 6) B. 3(4x + y + 10) C. 2(6x + 4y + 8) D. 2(6x + 5 + 2y)

ZombieLover:

So 2 × 6x = 12x, 2 × 5 = 10, 2 × 2y = 4y what doe sthat tell you?

supie:

So what you have to do is get the numbers outside of the parenthesis then distribute it to all the numbers inside of the parenthesis then if it equals 12x + 10 + 4y then thats the answer

darkknight:

12x + 10 + 4y? GCF is 2, undistribute the main one so you have 2( times what) = 12x+10+4y answer choice B is invalid, you cant take out a 3 and have whole numbers in the eqn